t...ex
Problem #3
Graph the inequality
F(x) > x
(Hint: rewrite in y=mx+b form and use the y-intercept and slope)
Problem # 17
A home-based company produces both hand-knitted scarves and sweaters. The scarves take 2 hours of labor to produce, and the sweaters take 14 hours. The labor available is limited to 40 hours per week, and the total production capacity is 5 items per week. Write a system of inequalities representing this situation, where x is the number of scarves and y is the number of sweaters. (You do not need to solve or graph.)

Answer by ankor@dixie-net.com(12687)

A home-based company produces both hand-knitted scarves and sweaters. The scarves take 2 hours of labor to produce, and the sweaters take 14 hours. The labor available is limited to 40 hours per week, and the total production capacity is 5 items per week. Write a system of inequalities representing this situation, where x is the number of scarves and y is the number of sweaters.
:
x = no. of scarves; y = no. of sweaters:
:
The labor inequality:
2x + 14y =< 40
:
Arrange in the y = mx + b, form:
14y =< 40 - 2x
y =< (40/14) - (2/14)x
y =< -(1/7)x + (20/7)
:
The production inequality:
x + y =< 5
y =< -x + 5
: